a capital of 500 dollars earns 150 dollars of interest in 6 years. What was the 

interest rate if compound interest is used? What if simple interest is used?


1
Expert's answer
2021-07-26T15:45:16-0400

Simple Interest ;

"150=\frac {500×6×r}{100}"


"r=\frac {150×100}{500×6}"


r=5%


Compound Interest ;

"C. I=P [(1 +\frac {r} {100} )^{n} - 1]"


"150=500[(1+\frac {r} {100})^6-1]"


"\frac {150}{500}+1=(1+\frac {r} {100})^6"


"\frac {650} {500}=(1+\frac {r} {100})^6"


"\frac {13}{10} =(1+\frac {r} {100})^6"


"(\frac {13}{10}) ^{\frac {1} {6}}=(1+\frac {r} {100})"


"r=[(\frac {13}{10}) ^{\frac {1} {6}} - 1]×100"


r=4.47%